# mmCooper: A Multi-agent Multi-stage Communication-efficient and Collaboration-robust Cooperative Perception Framework

### v1

@@ Abstract
Collaborative perception significantly enhances individual vehicle perception performance through the exchange of sensory information among agents. However, real-world deployment faces challenges due to bandwidth constraints and inevitable calibration errors during information exchange. To address these issues, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. Our framework leverages a multi-stage collaboration strategy that dynamically and adaptively balances intermediate- and late-stage information to share among agents, enhancing perceptual performance while maintaining communication efficiency. To support robust collaboration despite potential misalignments and calibration errors, our framework captures multi-scale contextual information for robust fusion in the intermediate stage and calibrates the received detection results to improve accuracy in the late stage. We validate the effectiveness of mmCooper through extensive experiments on real-world and simulated datasets. The results demonstrate the superiority of our proposed framework and the effectiveness of each component.
@@ 1 Introduction
With the development of autonomous driving systems and Internet of Vehicles (IoVs) technologies, perception devices such as cameras and LiDARs have been widely deployed. Although perception, as a critical module in autonomous vehicles, has witnessed rapid advancements driven by deep learning-based algorithms, the traditional single-vehicle perception paradigms [19, 24, 45] cannot meet the safety and reliability requirements of autonomous vehicles due to unavoidable factors such as object occlusion and limitations in detection range [53, 1, 12, 4, 22].
Figure 1: Comparison of different fusion approaches.
Benefiting from infrastructure improvements and the advancement of communication technologies like V2X [21, 20], autonomous vehicles can achieve multi-vehicle collaborative perception by sharing perception information [39, 3], which significantly enhances the performance of perception systems. Although cooperative perception has made progress recently, its practical deployment still faces challenges such as communication resource constraints [11, 6], localization errors [25, 26] and low information fusion efficiency [48, 38].
In cooperative perception systems, the effectiveness of fusion strategies significantly affects both perception performance and communication efficiency. As illustrated in Figure 1 (a)-(c), the fusion strategies can be divided into three categories: early, intermediate, and late fusion methods, illustrated. Specifically, early fusion methods [5, 53] aggregate raw sensor observations from all agents, providing abundant information that can significantly enhance perception performance. However, transmitting raw sensor observations requires substantial bandwidth, making this approach costly for communication. Intermediate fusion methods [11, 6, 48, 47, 46] address this issue by sharing encoded features rather than raw data to reduce communication overhead. While effective, feature compression inevitably introduces information loss, which can compromise fusion accuracy. Late fusion methods [30, 32] on the other hand, minimize communication demands by sharing lightweight perception results among agents. However, this approach is highly sensitive to environmental noise, as even minor disturbances can degrade collaborative quality. Overall, these above single-stage cooperative perception approaches face challenges in managing the trade-off between perception accuracy and communication efficiency, raising a compelling question: can we improve the perception performance while preserving communication efficiency by strategically sharing information across multiple stages to harness the strengths of each fusion stage?
@@
To this end, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. The proposed mmCooper framework facilitates effective information aggregation across fusion stages and enhances communication efficiency through carefully designed core components. Specifically, we design an adaptive multi-stage data fusion mechanism that facilitates multi-agent cooperation by dynamically fusing information at both intermediate and late stages, as shown in Figure 1. Guided by a confidence-based filtering gaile strategy, this mechanism automatically assesses and determines the data volume to be processed at intermediate and late stages. High-confidence bounding boxes are shared directly, while intermediate features are selectively shared instead of bounding boxes for regions requiring further contextual reference, and low-confidence areas are excluded to prevent propagating misleading information. This dual-stage fusion approach enhances perception performance while maintaining communication efficiency.
Besides, to further address potential misalignment and calibration noise among agents, we incorporate specific designs in the proposed mmCooper framework for both intermediate and late stages. In the intermediate stage, we introduce a Multi-scale Offset-aware Fusion Module that not only fuses data from target locations related to cooperative agents’ views but also from nearby regions, adding redundancy to mitigate calibration noise. For the late stage, we design a Bounding Box Filtering & Calibration Module that uses the information-rich intermediate-stage features from the ego agent to filter inaccurate bounding boxes received from other agents and refine the remaining noisy ones. By integrating these corrected and ego-generated bounding boxes, our mmCooper framework is able to achieve high perception accuracy and robustness. We evaluated the performance of our mmCooper framework on both real and virtual datasets, including DAIR-V2X [49], OPV2V [43] and V2XSet[42]. Experimental results demonstrate that mmCooper outperforms previous methods across various settings, confirming its effectiveness as a communication-efficient and robust multi-stage cooperative perception framework. To summarize, our main contributions are summarized as follows:
@@
  • To address the potential misalignment and calibration error among agents, we incorporate a Multi-scale Offset-aware Fusion Module to integrate spatially adjacent contextual information at different feature scales in the intermediate stage and a Bounding Box Filtering & Calibration Module to filter and improve the bounding boxes from other agents in the late stage, guided by the information-rich intermediate features.
  • Extensive experiments show excellent performance on both real-world and simulated datasets, with a 7.29%/1.31%/2.09% improvement in AP@0.7 on OPV2V, DAIR-V2X and V2XSet datasets over the second-best state-of-the-art methods. Additionally, our approach requires only 1/9153, 1/156 and 1/18305 of the communication volume used by comparable methods.
@@ 2 Related Work
Figure 2: Overview of mmCooper framework: a cooperative perception system with adaptive multi-stage fusion.
@@ 2.2 Collaborative Perception
Collaborative perception addresses the limitations of single-agent perception by aggregating perception information from surrounding agents to adapt to complex and dynamic autonomous driving environments. Current work is categorized into early, intermediate, and late fusion based on collaboration stages. Early fusion [5, 51, 53] shares raw point cloud data among agents, achieving high perception performance but consuming significant bandwidth resources. EMP [53] achieved scalability and adaptability on highly fluctuating open roads by dynamically partitioning the point cloud range shared by agents. Late fusion [30, 44, 32] shares lightweight proposals between agents but performs poorly and is sensitive to noise. VIPS [30] implemented efficient graph-structured matching to enable object-level fusion. Intermediate fusion [23, 11, 42, 50, 18] shared features among agents, yet complete Intermediate features still incur high bandwidth consumption. Recent work focused on novel communication mechanisms to filter Intermediate features and reduce communication volume. for instance, Where2comm [11] employed a confidence-based scheme to guide agents in focusing on sharing spatially critical information. ERMVP [50] adopted a hierarchical feature sampling strategy to reduce communication overhead while leveraging sparse consensus features to mitigate localization errors. Different from these works, our framework conducts multi-stage fusion instead of single-stage which balances the perception performance and communication overhead. As far as we know, only a few works have considered multi-stage settings. Zhang et al. [52] proposes a defense method that combines raw data and intermediate features to fuse redundant information, thereby reducing the impact of spurious data. Zhao et al. [54] simultaneously considers early raw data and late-stage detection results to address malicious agents in collaborative environments. However, both works focus on security issues related to collaborative perception instead of improving sensing performance. Xie et al. [40] applies reinforcement learning for dynamic partitioning of early, intermediate, and late-stage information. However, this work merely divides the data without integrating information across stages. In contrast to these works, this paper introduces a collaborative perception framework that aims to improve sensing performance by conducting multi-stage data fusion.
@@ 3.1 Overview
This section introduces our proposed mmCoop framework, as illustrated in Figure 2, which comprises four main components: the Observation Module, the Broadcasting Module, the Intermediate-stage Fusion, and the Late-stage Fusion. The Observation Module (Sec.3.2) serves as the foundation by extracting essential features from the ego vehicle’s LiDAR-based point cloud data. The Broadcasting Module (Sec.3.3) includes a Detection Decoder to generate coarse bounding boxes from intermediate features, along with a Confidence-based Filter Generation Module to produce filters that dynamically determine the data volume shared at the intermediate and late stages. These generated filters enable adaptive data sharing: high-confidence bounding boxes are shared directly, intermediate features are selectively transmitted for areas needing additional context, and low-confidence areas are excluded, balancing perception accuracy with communication efficiency. The Intermediate-stage Fusion (Sec.3.4) contains the Multi-scale Offset-aware Fusion Module, which uses cross-attention to integrate features from both target and nearby regions for robust feature fusion. In the Late-stage Fusion (Sec.3.5), a Bounding Box Filtering and Calibration Module filters inaccurate bounding boxes from other agents and calibrates the remaining ones using the ego agent’s intermediate features. The Late-stage Fusion contains the Bounding Box Filtering and Calibration Module to filter out inaccurate bounding boxes received from other agents and refine the remaining noisy ones by leveraging the information-rich intermediate features from the ego agent. Then, the calibrated bounding boxes from all other agents will be combined with the bounding boxes generated by the ego agent to produce the final prediction.
3.2 Observation Module
@@
In a multi-agent cooperative perception scenario, consider the collaboration of $n$ agents, represented by the agent set $N=\left\{1,...,n\right\}$. One vehicle serves as the ego vehicle and the remaining $n-1$ agents act as collaborators. The ego vehicle aggregates information from the collaborators to achieve cooperative perception. Considering the i-th agent as the ego vehicle, for the $j$-th collaborator, where $i,j\in N$, the point cloud observation is given as $X_{j}$. The ego vehicle first broadcasts its basic information to the collaborators, such as coordinates and heading angle. The collaborators utilize this basic information to project their point cloud observations into the ego vehicle’s coordinate system, with the projected point cloud observation represented as $x_{j}^{{}^{\prime}}=T_{j->i}\cdot x_{j}$, where $T_{j->i}$ denotes the transformation matrix from agent j to agent i. Subsequently, each agent extracts features from its point cloud observation through a shared encoder to obtain Bird’s Eye View (BEV) features, represented as $F_{bev}^{k}=\psi_{encoder}\left(x_{k}\right)\in\mathbb{R}^{C\times H\times W}$, where $\psi_{encoder}$ denotes the shared PointPillar encoder [15], and $k\in N$, and $C$, $H$, and $W$ correspond to the channel, height, and width of the feature map, respectively.
3.3 Broadcasting Module
Figure 3: The Confidence-based Filtering Generation Module
For the locally encoded BEV features, prior work has mainly focused on either directly sharing intermediate-stage features or decoding them to obtain late-stage bounding boxes, which are then shared. However, collaboration schemes based on a single stage often struggle to effectively balance communication bandwidth and perception performance. To effectively leverage the advantages of multi-stage data, we introduce a multi-stage collaboration framework that simultaneously balances bandwidth and accuracy by transferring filtered features and proposals between agents. As shown in Figure 3, our Broadcasting Module filters the intermediate features and initially generates coarse bounding boxes, which are then transmitted to other agents.
Confidence-based Filter Generation Module. The Confidence-based Filter Generation Module, as a key component of the Broadcasting Module, serves two main purposes: 1) dynamically and adaptively determining the allocation of features and coarse bounding boxes; 2) ensuring that the discrete operations involved in obtaining the filter can still propagate gradients properly. To filter spatial features and bounding boxes, we input the encoded features into the Confidence Generator to generate spatial confidence maps for both the features and proposals. The confidence maps represent the potential regions of objects and are denoted as:
$C_{f}^{i},C_{b}^{i}=\phi_{conf}\left(F_{bev}^{i}\right),$
where $\phi_{conf}\left(\cdot\right)$ refers to the Multi-stage Confidence Generator, and $C_{f}^{i},C_{b}^{i}$ represents the confidence maps for the features and proposals, respectively. A Gaussian Filter is then applied to the spatial confidence maps, helping to filter out anomalous areas and incorporate contextual effects, thereby enhancing the robustness of critical region selection. To eliminate irrelevant background information, the top $m\%$ highest-confidence positions are selected from each confidence map to generate the spatial filter $M_{f/b,top}^{i}$. Then we add Gumbel noise $g_{f/b}$ to the confidence maps to introduce a certain level of randomness, expressed as:
$g_{f/b}=-log\left(-logu\right),u\sim Uniform\left(0,1\right).$
The introduction of Gumbel noise can enhance the exploratory nature of the model and improve its generalization capability. For the same positions in the confidence maps from two stages, we retain only the parts with relatively high confidence scores. The confidence score indicates the potential location of an object. When a high-dimensional bounding box exhibits high confidence in representing the object, the agent will share the bounding box. Otherwise, intermediate-stage features are required to convey the information at that location. In addition, to ensure the continuity of the dynamic allocation process for the intermediate-stage features and coarse bounding boxes. we employ the Gumbel Softmax method to obtain approximate samples from a discrete distribution. Utilizing the straight-through estimator to acquire discrete distributions while preserving standard forward propagation. The specific process is represented as follows:
$M_{f/b,g}^{i}=f_{max}\left(\frac{exp\left((logC_{f/b}+g_{f})/\tau\right)}{{{\textstyle\sum_{s}}exp\left((logC_{s}+g_{s})/\tau\right)}}\right),$
@@ 3.3 Broadcasting Module
where $\tau$ denotes the temperature parameter of the distribution $(\tau>0)$, and $s\in{\left\{f,b\right\}}$. Based on the above process, our final filter $\mathcal{M}_{f/b}^{i}$ is represented as:
$\mathcal{M}_{f/b}^{i}=M_{f/b,g}^{i}\odot M_{f/b,top}^{i},$

### v3

@@ Abstract
Collaborative perception significantly enhances individual vehicle perception performance through the exchange of sensory information among agents. However, real-world deployment faces challenges due to bandwidth constraints and inevitable calibration errors during information exchange. To address these issues, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. Our framework leverages a multi-stage collaboration strategy that dynamically and adaptively balances intermediate- and late-stage information to share among agents, enhancing perceptual performance while maintaining communication efficiency. To support robust collaboration despite potential misalignments and calibration errors, our framework prevents misleading low-confidence sensing information from transmission and refines the received detection results from collaborators to improve accuracy. The extensive evaluation results on both real-world and simulated datasets demonstrate the effectiveness of the mmCooper framework and its components.
@@ 1 Introduction
With the advancement of autonomous driving systems [58, 19], perception devices such as cameras and LiDARs have been widely deployed. Although perception technologies have witnessed rapid advancements driven by deep learning-based algorithms, the traditional single-vehicle perception paradigms [23, 30, 53] cannot meet the safety and reliability requirements of autonomous vehicles due to unavoidable factors such as object occlusion and limitations in detection range [62, 1, 14, 6, 27].
Benefiting from infrastructure improvements and the advancement of Internet of Vehicles (IoVs) technologies like V2X [26, 25], autonomous vehicles can achieve multi-vehicle collaborative perception by sharing perception information [47, 5], which significantly enhances the performance of perception systems. Although cooperative perception has made progress recently, its practical deployment still faces challenges such as communication resource constraints [13, 8], localization errors [31, 34], and low information fusion efficiency [56, 46].
In cooperative perception systems, the effectiveness of fusion strategies significantly affects both perception performance and communication efficiency. The fusion strategies can be divided into three categories: early, intermediate, and late fusion methods. Specifically, early fusion methods [7, 62] aggregate raw sensor observations from all agents, providing abundant information but requiring immense bandwidth, making this approach costly for communication. Intermediate fusion methods [13, 8, 56, 55, 54] partially mitigate this issue by sharing encoded features rather than raw data to reduce communication overhead. However, transmitting the feature map of the entire sensing scenario still results in substantial bandwidth consumption [13]. Late fusion methods [38, 40] on the other hand, minimize communication demands by sharing lightweight perception results among agents. However, this approach is highly sensitive to environmental noise, as even minor disturbances can degrade collaborative quality [31]. Overall, these above single-stage cooperative perception approaches face challenges in managing the trade-off between perception accuracy and communication efficiency, raising a compelling question: Can we improve the perception performance while preserving communication efficiency by strategically sharing information across multiple stages to harness the strengths of each fusion stage?
@@
To this end, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. The proposed mmCooper framework facilitates effective information aggregation across fusion stages and enhances communication efficiency through carefully designed core components. Specifically, we design an adaptive multi-stage data fusion mechanism that facilitates multi-agent cooperation by dynamically fusing information at both intermediate and late stages. Guided by a confidence-based filtering strategy, this mechanism automatically assesses and determines the data volume to be processed at intermediate and late stages. High-confidence bounding boxes are shared directly, while intermediate features are selectively shared instead of bounding boxes for regions requiring further contextual reference, and low-confidence areas are excluded to prevent propagating misleading information. This dual-stage fusion approach enhances perception performance while maintaining communication efficiency.
Besides, to further address potential misalignment and calibration noise among agents, we incorporate specific designs in the proposed mmCooper framework for both intermediate and late stages. In the intermediate stage, we introduce a Multi-scale Offset-aware Fusion Module that not only fuses data from target locations related to cooperative agents’ views but also from nearby regions, adding redundancy to mitigate calibration noise. For the late stage, we design a Bounding Box Filtering & Calibration Module that uses the information-rich intermediate-stage features from the ego agent to filter inaccurate bounding boxes received from other agents and refine the remaining noisy ones. By integrating these corrected and ego-generated bounding boxes, our mmCooper framework can achieve high perception accuracy and robustness. To summarize, our main contributions are summarized as follows:
@@
  • To address the potential misalignment and calibration error among agents, we design a Multi-scale Offset-aware Fusion Module to integrate spatially adjacent contextual information at different feature scales in the intermediate stage and a Bounding Box Filtering & Calibration Module to filter and improve the bounding boxes from other agents in the late stage, guided by the information-rich intermediate features.
  • Extensive experiments show excellent performance on both real-world and simulated datasets, with a 7.29%/1.31%/2.09% improvement in AP@0.7 on OPV2V [50], DAIR-V2X [57] and V2XSet [49] datasets over the second-best SOTA methods. Meanwhile, our approach requires only 1/9153, 1/156, and 1/18305 of the communication volume used by comparable methods. Furthermore, our method consistently outperforms previous approaches under various configurations.
@@ 2.2 Collaborative Perception
Collaborative perception addresses the limitations of single-agent perception by aggregating perception information from surrounding agents to adapt to complex and dynamic autonomous driving environments. Current work is categorized into early, intermediate, and late fusion based on collaboration stages. Early fusion [7, 60, 62] shares raw point cloud data among agents, achieving high perception performance but consuming significant bandwidth resources. EMP [62] achieved scalability and adaptability on highly fluctuating open roads by dynamically partitioning the point cloud range shared by agents. Late fusion [38, 51, 40] shares lightweight proposals between agents but performs poorly and is sensitive to noise. VIPS [38] implemented efficient graph-structured matching to enable object-level fusion. Intermediate fusion [29, 13, 49, 59, 22] shared features among agents, yet complete Intermediate features still incur high bandwidth consumption. Recent work focused on novel communication mechanisms to filter Intermediate features and reduce communication volume. for instance, Where2comm [13] employed a confidence-based scheme to guide agents in focusing on sharing spatially critical information. ERMVP [59] adopted a hierarchical feature sampling strategy to reduce communication overhead while leveraging sparse consensus features to mitigate localization errors. Different from these works, our framework conducts multi-stage fusion instead of single-stage which balances the perception performance and communication overhead. As far as we know, only a few works have considered multi-stage settings. Zhang et al. [61] proposes a defense method that combines raw data and intermediate features to fuse redundant information, thereby reducing the impact of spurious data. However, this work focus on security issues related to collaborative perception instead of improving sensing performance. Xie et al. [48] applies reinforcement learning for dynamic partitioning of early, intermediate, and late-stage information. However, this work merely divides the data without integrating information across stages. In contrast to these works, this paper introduces a collaborative perception framework that aims to improve sensing performance by conducting multi-stage data fusion.
Figure 1: Comparison between our proposed multi-stage mmCooper framework and existing methods. The BEV (Bird’s Eye View) communication map illustrates the information shared by agents at each scene location. (a)(b)(c) depict existing methods, which transmit the entire scene’s data in a single stage without considering sensing confidence, leading to excessive communication overhead and degraded performance. (d) demonstrates how mmCooper selectively transmits non-overlapping information across multiple stages, reducing communication costs and enhancing model performance.
@@ 3.1 Overview
As illustrated in Fig. 1, our key idea is to conduct selective multi-stage agent collaboration, rather than the single-stage fusion used in existing methods. Conventional approaches indiscriminately transmit data encompassing the entire scene through a single stage, introducing undesired low-quality sensing information during communication. This not only increases communication overhead but also complicates fusion, as the system must first evaluate the quality of received data to ensure reliable results. In contrast, our approach adopts a selective multi-stage collaboration strategy, where the transmission decision, whether to send data and at which stage, is determined based on the sensing confidence level of each location. For high-confidence regions, we transmit late-stage bounding boxes, minimizing communication costs. For moderate-confidence regions, we transmit intermediate-stage features for richer information. For low-confidence regions, transmission is suppressed to prevent misleading other agents. Additionally, our framework incorporates alignment and calibration corrections during both intermediate and late-stage fusion, further enhancing the reliability of the fused results.
@@
The details of our proposed mmCooper framework are illustrated in Fig. 2, which comprises three main components: (1) the Information Broadcasting, (2) the Intermediate-stage Fusion, and (3) the Late-stage Fusion. The encoded point cloud features will be sent to two branches. In one branch, the Information Broadcasting (Sec. 3.3) employs a Confidence-based Filter Generation Module to dynamically determine whether data should be suppressed, transmitted at the intermediate stage, or transmitted at the late stage. In the other branch, the Intermediate-stage Fusion (Sec. 3.4) contains the Multi-scale Offset-aware Fusion Module, which uses cross-attention to integrate features from both target and neighboring regions. This module mitigates potential misalignment between features from the ego and other agents, ensuring robust feature fusion. In Late-stage Fusion (Sec. 3.5), a BBox Filtering & Calibration Module removes inaccurate bounding boxes from other agents and refines the remaining ones using the ego agent’s information-rich intermediate features. The calibrated bounding boxes from all agents will be merged with those generated by the ego agent to produce the final outputs.
Figure 2: Overview of mmCooper framework: a cooperative perception system with adaptive multi-stage fusion. It consists of three main components: (1) Information Broadcasting (Sec. 3.3) filters features and bounding boxes to achieve bandwidth efficiency; (2) Intermediate-stage Fusion (Sec. 3.4) captures surrounding information from the received feature maps for robust feature fusion; (3) Late-stage Fusion (Sec. 3.5) utilizes the information-rich fused features for Filtering & Calibration of the received bounding boxes.
3.2 Observation Encoding
@@
In a multi-agent cooperative perception scenario, consider the collaboration of $n$ agents, represented by the agent set $N=\left\{1,...,n\right\}$. The $i$-th agent serves as the ego vehicle and the remaining n1n-1 agents act as collaborators. All the agents first broadcast basic positional information to their collaborators, including coordinates and heading angle. In this way, the agents obtain the necessary information to project the information from the collaborators’ systems to the ego vehicle’s coordinate system. The ego agent processes its point cloud through a shared encoder to extract Bird’s Eye View (BEV) features, represented as $F^{i}=\psi_{E}\left(x^{i}\right)\in\mathbb{R}^{C\times H\times W}$, where $\psi_{E}$ denotes the shared PointPillar [18] encoder, $x^{i}$ represents the projected local observation, and $C$, $H$, and $W$ correspond to the channel, height, and width of the feature map, respectively.
3.3 Information Broadcasting
As shown in the top-left corner of Fig. 2, when conducting information broadcasting, the Confidence-based Filter Generation (CFG) Module dynamically decides whether data at each location should be suppressed, transmitted as intermediate-stage features, or transmitted as late-stage bounding boxes. Notably, no additional model or data compression techniques [32] are applied in this framework, only raw features or bounding boxes are transmitted. The bounding boxes are generated by the Detection Decoder from the features and the transmitting decision is guided by the confidence filter generated by the Filter Generator (Fig. 3) based on the sensing confidence, as detailed below.
Figure 3: The design of Filter Generator in CFG Module, generating confidence scores to guide information transmission at each location.
Filter Generator. The encoded BEV features (i.e., $F^{i}$) are fed into this Confidence Generator to produce intermediate-stage confidence map $C_{f}^{i}$ and late-stage confidence map $C_{b}^{i}$ for all the locations of the scene using CNN. A Gaussian Filter is then applied to the generated confidence maps to smooth out noise and reduce anomalies, improving the selection of critical regions. To suppress the information of the locations with low confidence, only the top $p\%$ highest-confidence positions are selected from each confidence map to generate the spatial filters $M_{f/b,t}^{i}$ (i.e., $M_{f,t}^{i}$ or $M_{b,t}^{i}$). Then, for each location, a classification is required to decide which stage to transmit the data. However, using the traditional Softmax will prevent the back-propagation of the gradient due to its non-differentiability as a discrete operation. Instead, we utilize the Gumbel Softmax [16] to obtain approximate samples from a discrete distribution ensuring the propagation of gradient through the straight-through estimator [2] while preserving standard forward propagation. The Gumbel filter $M_{f/b,g}^{i}$ obtained in this process is given by the following equation:
$M_{f/b,g}^{i}=f_{max}\left(\frac{exp\left((logC_{f/b}^{i}+g_{f/b})/\tau\right)}{{{\textstyle\sum_{s}}exp\left((logC_{s}^{i}+g_{s})/\tau\right)}}\right),$
@@ 3.3 Information Broadcasting
where $g_{f/b}$ is Gumbel noise [33], $\tau$ denotes the temperature parameter of the distribution $(\tau>0)$, and $s\in{\left\{f,b\right\}}$. Based on the above process, our final filter $\mathcal{M}_{f/b}^{i}$ is represented as:
$\mathcal{M}_{f/b}^{i}=M_{f/b,g}^{i}\odot M_{f/b,t}^{i}.$
Based on the aforementioned filter, the intermediate-stage feature $F^{i}$ and the coarse bounding box $\mathcal{B}^{i}$ are filtered to obtain the filted feature $\hat{F}^{i}$ and coarse bounding box $\hat{\mathcal{B}}^{i}$.